2/3x=8/x+6-2/3 you want to get your x's on the same side so I will take the minus 2/3 and add it to the 2/3x which will cancel it out and leave you with x=8/x+6 then I want to get the 8/x on the side with the other x so I will subtract it and get -7x=6 the I will divide the -7 by 6 and will get -17/20 and your answer will be x=-17/20
Answer by alytoland(1)

2/3x=8/x+6-2/3 you want to get your x's on the same side so I will take the minus 2/3 and add it to the 2/3x which will cancel it out and leave you with x=8/x+6 then I want to get the 8/x on the side with the other x so I will subtract it and get -7x=6 the I will divide the -7 by 6 and will get -17/20 and your answer will be x=-17/20